OpenStudy (anonymous):

What are the zeros of the polynomial function: f(x) = x^3 – 5x^2 + 6x

OpenStudy (cwrw238):

first take out x: f(x) = x(x^2 - 5x + 6) = 0 (x(x - 3)(x - 2) = 0 can you continue?

OpenStudy (anonymous):

i still do no understand

OpenStudy (anonymous):

not*

OpenStudy (cwrw238):

did you understand why i took the x out?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

0, –3, 2 0, 3, –2 0, –3, –2 0, 3, 2 I got these four answers

OpenStudy (anonymous):

do you understand why there's three answers?

OpenStudy (anonymous):

yes but i dont know which one is correct, i tried four different times and ended up with these ugh

OpenStudy (anonymous):

The zeroes is just where y = 0. So pretend f(x) is your y and plug in 0. 0 = x^3 – 5x^2 + 6x Do you know how to solve this? Just ignore f(x)

OpenStudy (anonymous):

I would start by factoring first, all of these numbers have a common factor, x. x (x^2 - 5x + 6) Now, factor it further by factoring x^2-5x+6 You get (x-3)(x-2) you're left with x(x-3)(x-2) Equal each of these expressions to zero. x=0 x - 3 = 0, which is x=3 x - 2 = 0, which is x =2